# Cohomology of Lie algebroids over Algebraic spaces

### v7

@@ Abstract.
AMS Mathematics Subject Classification : $16$S30303030, $16$T10,1010,10 ,$17$B56565656.
@@ 1. Introduction
The notion of Lie algebroids plays a prominent role in differential geometry and mathematical physics as they represent generalized infinitesimal symmetries of spaces, which are related to the corresponding global symmetries of spaces described by Lie groupoids [21]. Lie algebroids over a $C^{\infty}$-manifold are joint generalization of tangent vector bundle over the manifold and Lie algebras [26]. An algebraic analogue of Lie algebroids, known as Lie-Rinehart algebras, is used to study general situations [16, 18, 22, 3]. In the context of complex geometry and algebraic geometry, Lie algebroids over analytic spaces [32] and over algebraic varieties (or schemes) [18, 7, 3] respectively have been studied in sheaf theoretic language, which are joint generalization of the tangent sheaf and sheaves of Lie algebras [18, 33, 32]. We study Lie algebroids over an algebraic space (shortly, $a$-space) or over a topological ringed space $(X,\mathcal{O}_{X})$ as certain quasicoherent sheaves of Lie-Rinehart algebras [37, 29], where $\mathcal{O}_{X}$ is a sub-sheaf of algebras of the sheaf of continuous functions $C_{X}^{0}$ on $X$ [33, 30]. This framework unifies the concept of Lie algebroids across these three types of base spaces as special cases. To study calculus on these (smooth and singular) geometric objects in a consistent manner, we need the notion of Lie algebroids within algebro-geometric settings. It allows one to treat several geometric structures, such as Poisson analytic spaces [32], singular foliations or generalized involutive distributions [10, 32, 20], as well as free Lie algebroids [18]. In particular, the sheaf of logarithmic derivations for some (principal or free) divisor of a complex manifold (or a smooth algebraic variety) [4, 32, 29], and the path algebroid of a smooth manifold [18], are key object of study in this context.
@@
In Section 2, we recall some of the preliminary notions required for the article. In Section 3, we consider Lie algebroid hypercohomology over certain special $a$-spaces. Here, we show that one can define Lie algebroid cohomology, where the underlying $\mathcal{O}_{X}$-module, while locally free, is not necessarily of finite rank, thus encompassing all three geometric setups. The cohomology of a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with coefficient in a $\mathcal{L}$-module $\mathcal{E}$, is given by $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$. Here, we use the result of Lie-Rinehart cohomology as derived functor for a projective Lie-Rinehart algebra [34, 16] as a local description. We apply this result for the sheaf of logarithmic derivations and for $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras. Using the cohomology groups we consider (first) Chern classes [24, 25] in this general set up. Furthermore, we express the Lie algebroid cohomology in terms of Čech cohomology by considering a good open cover [33, 36, 1]. In Section 4, we consider algebraic (analytic) de Rham cohomologies for some free divisors associated with principal ideal sheaves [29] and compute the corresponding logarithmic de Rham cohomologies [4, 27]. We view these cohomologies as Lie algebroid (hyper)cohomology as described in [32, 1, 3].
@@ 2.1. Lie algebroids over algebraic spaces
For an $a$-space $(X,\mathcal{O}_{X})$, the sheaf of derivations $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$ of $\mathcal{O}_{X}$, known as the tangent sheaf of $X$.
@@
Consider smooth manifolds, complex manifolds, analytic spaces and algebraic varieties, all with their associated structure sheaf, as special $a$-spaces. The tangent sheaf over such an $a$-space $(X,\mathcal{O}_{X})$ is a quasicoherent module over $\mathcal{O}_{X}$ [18, 32].
@@
A homomorphism of Lie algebroids over an $a$-space
@@
is a sheaf homomorphism of Lie-Rinehart algebras.
@@
The standard Lie algebroid structure on the tangent sheaf of an $a$-space $(X,\mathcal{O}_{X})$ is described by the pair $(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c},id)=:\mathcal{T}_{X}$. In particular, when $X$ is a real smooth manifold (complex manifold), the sheaf of smooth (holomorphic) vector fields $\mathfrak{X}_{X}$ on $X$ is isomorphic to $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, forming a locally free Lie algebroid of rank equals to dimension of the manifold $X$.
@@ 2.2. Universal enveloping algebroid of a Lie algebroid
From the construction of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, it is an associative $\mathbb{K}_{X}$-algebra and $\mathcal{O}_{X}$-bimodule. Moreover, there is a canonical $\mathbb{K}_{X}$-algebra monomorphism $\iota_{X}:\mathcal{O}_{X}\hookrightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and an $\mathcal{O}_{X}$-linear map $\iota_{\mathcal{L}}:\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. The sheaf $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is generated by $\mathcal{O}_{X}$ and $\iota_{\mathcal{L}}(\mathcal{L})$ satisfy the following identities:
@@ 2.3. $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras
  • $\mathcal{A}$ is equipped with the morphism of sheaf of $\mathcal{O}_{X}$-modules the comultiplication $\Delta:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathcal{O}_{X}}\mathcal{A}$ and counit $\epsilon:\mathcal{A}\rightarrow\mathcal{O}_{X}$;
@@
For a smooth ((((holomorphic)))) Lie algebroid [1] i.e. for a locally free sheaf of Lie-Rinehart algebras of finite rank over a smooth (complex) manifold, the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of finite type)))) [29].
The tangent sheaf over the affine scheme $Spec(\mathbb{A}^{\mathbb{N}})$ ((((where $\mathbb{A}^{\mathbb{N}}:=\mathbb{K}[x_{i}]_{i\in\mathbb{N}})$ [9, 29] and the free Lie algebroid $\mathcal{P}_{X}$ over a smooth manifold $X$ [18] are locally free sheaf of Lie-Rinehart algebras ((((of infinite rank)))). Thus, their universal enveloping algebroids are cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras ((((of infinite type)))) [29].
@@ 3. Lie algebroid Cohomology over $a$-spaces
Lie algebroids over algebraic spaces (shortly, $a$-spaces) are discussed in Section 2.1 by considering certain quasicoherent sheaves of Lie-Rinehart algebras over $a$-spaces $(X,\mathcal{O}_{X})$ (form an abelian category). The category of representations of a Lie algebroid has enough injectives (since the category of quasicoherent sheaves of left $\mathcal{O}_{X}$-modules is enough injectives [14]), useful for doing homological algebra.
@@ 3.1. Atiyah algebroid
In particular, when $\mathcal{E}=\mathcal{O}_{X}$ for a non singular $a$-space $X$, we have $\mathcal{A}t(\mathcal{O}_{X})\cong\mathcal{O}_{X}\oplus\mathcal{T}_{X}$, equals to the sheaf differential operators of order $\leq 1$ with scalar symbol. The universal enveloping algebroid of the Atiyah algebroid $\mathcal{A}t(\mathcal{O}_{X})$ is the sheaf of twisted differential operators over $X$ [2].
@@
For a $\mathcal{O}$-module $\mathcal{C}$, we can canonically form the following sheaves of graded vector spaces $\wedge^{\bullet}_{\mathcal{O}}$ $\mathcal{C}:=\oplus_{i}\wedge^{i}_{\mathcal{O}}\mathcal{C}$ and $\mathcal{C}^{\otimes^{\bullet}}:=\oplus_{i}$ $\mathcal{C}^{\otimes^{i}}$ (forms exterior algebra and tensor algebra with respect to the products $\wedge$ and $\otimes$ over $\mathcal{O}$ respectively).
@@
A quasi-isomorphism of complexes $f^{\bullet}:\mathcal{F}^{\bullet}\rightarrow\mathcal{G}^{\bullet}$ induces an isomorphism $\mathbb{H}^{k}(X,f^{\bullet})$, and if $\mathcal{I}^{\bullet}$ is a complex of injective sheaves then $\mathbb{H}^{k}(X,\mathcal{I}^{\bullet})=H^{k}(\Gamma(X,\mathcal{I}^{\bullet}))$.
@@
We use sheaf-theoretic generalizations of the well-known derived functors $Ext$ and $Cotor$, by sheafifying the standard cochain complexes ((((see [34], [19])))) and subsequently considering the associated hypercohomologies.
@@ 3.2. Chevalley-Eilenberg-de Rham complex and Koszul-Rinehart resolution
where $D_{1},\dots,D_{k+1}\in\mathcal{L}(U)$ and $\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U))$, and $\nabla^{U}:=\nabla(U)$ (see the notion in (2), (3)). Notice that, the differential $d$ (i.e. $d^{2}=0)$ is a $\mathbb{K}_{X}$-linear map, but not an $\mathcal{O}_{X}$-module homomorphism ((((satisfies the Leibniz rule: $d(f~{}\omega)=df\wedge\omega+f~{}d\omega$, for $f\in\mathcal{O}_{X}(U),\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U)))$.
@@
which implies $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\Omega^{\bullet}_{\mathcal{L}}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ where $\Omega^{\bullet}_{\mathcal{L}}:=\Omega^{\bullet}_{\mathcal{L}}(\mathcal{O}_{X})$. This complex is equivalent to the standard complex associated with a smooth Lie algebroid [21] or a holomorphic Lie algebroid [32]. In particular, we get the Chevalley-Eilenberg complex when $\mathcal{L}$ is a Lie algebroid over a point; and the de Rham complex $\Omega^{\bullet}_{X}$ when $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}=\mathcal{O}_{X}$ over a non-singular (or smooth) $a$-space $X$.
@@
Consider the chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules (call it by Koszul-Rinehart complex)
@@
it is the sheafification of the presheaf of cochain complexes [34, 39]
@@
For every $x\in X$, the space of sections $\mathcal{L}(U_{x})$ is a $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra and a free module over $\mathcal{O}_{X}(U_{x})$. Thus, the Koszul-Rinehart resolution (7) is a locally free resolution of the sheaf of local ring $\mathcal{O}_{X}$.
Here, we consider the two naturally associated presheaf of cochain complexes
@@
One can find the proof for a locally free Lie algebroid $\mathcal{L}$ of finite rank over a Noetherian separated scheme $(X,\mathcal{O}_{X})$ discussed in [3], by using the ideas of [34] on the level of stalks $\mathcal{L}_{x}$ (since $\mathcal{L}_{x}$ is a $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebra, projective (in fact, free) module over $\mathcal{O}_{X,x}$ for all $x\in X)$.
@@
Consider the path algebroid $\mathcal{P}_{X}$ for a non-singular special $a$-space $(X,\mathcal{O}_{X})$ and its universal enveloping algebroid $\mathbb{D}_{X}$ (see Remark 2.12). Note that a module over the free Lie algebroid $\mathcal{P}_{X}$ can be viewed as a vector bundle with connection and a $\mathbb{D}_{X}$-module describe some system of PDE on the space of paths (see [18] for details).
@@
Also, note that the whole ideas works for the classical case of Lie algebroids over smooth manifolds (i.e. for smooth Lie algebroids). In [1], the Lie algebroid ((((hyper))))cohomology of holomorphic ((((algebraic)))) Lie algebroids over a complex manifold ((((smooth scheme)))) has expressed by Čech cohomology using a good open cover consisting with connected Stein manifolds ((((affine schemes)))). However, the case associated with analytic spaces is not considered. Additionally, this result can be extended to quasicoherent $\mathcal{O}_{X}$-modules.

### v8

@@ Abstract.
AMS Mathematics Subject Classification : $16$S30303030, $16$T10,1010,10 ,$17$B56565656, $32$L10101010, $53$D17171717.
@@ 1. Introduction
The notion of Lie algebroids plays a prominent role in differential geometry and mathematical physics as they represent generalized infinitesimal symmetries of spaces, which are related to the corresponding global symmetries of spaces described by Lie groupoids [21]. Lie algebroids over a $C^{\infty}$-manifold are joint generalization of tangent vector bundle over the manifold and Lie algebras [26]. An algebraic analogue of Lie algebroids, known as Lie-Rinehart algebras, is used to study general situations [16, 18, 22, 3]. In the context of complex geometry and algebraic geometry, Lie algebroids over analytic spaces [32] and over algebraic varieties (or schemes) [18, 7, 3] respectively have been studied in sheaf theoretic language, which are joint generalization of the tangent sheaf and sheaves of Lie algebras [18, 33, 32]. We study Lie algebroids over an algebraic space (shortly, $a$-space) or over a topological ringed space $(X,\mathcal{O}_{X})$ as certain quasicoherent sheaves of Lie-Rinehart algebras [18, 38, 29], where $\mathcal{O}_{X}$ is a sub-sheaf of algebras of the sheaf of continuous functions $C_{X}^{0}$ on $X$ [33, 30]. This framework unifies the concept of Lie algebroids across these three types of base spaces as special cases. To study calculus on these (smooth and singular) geometric objects in a consistent manner, we need the notion of Lie algebroids within algebro-geometric settings. It allows one to treat several geometric structures, such as Poisson analytic spaces [32], singular foliations or generalized involutive distributions [10, 32, 20], as well as free Lie algebroids [18]. In particular, the sheaf of logarithmic derivations for some (principal or free) divisor of a complex manifold (or a smooth algebraic variety) [4, 32, 29], and the path algebroid of a smooth manifold [18], are key object of study in this context.
@@
In Section 2, we recall some of the preliminary notions associated with a Lie algebroid that are required for the article. In Section 3, we consider Lie algebroid hypercohomology over certain special $a$-spaces. Here, we show that one can define Lie algebroid cohomology, where the underlying $\mathcal{O}_{X}$-module, while locally free, is not necessarily of finite rank, thus encompassing all three geometric setups. The cohomology of a locally free Lie algebroid $\mathcal{L}$ over an $a$-space $(X,\mathcal{O}_{X})$ with coefficient in a $\mathcal{L}$-module $\mathcal{E}$, is given by $\mathbb{H}^{\bullet}(\mathcal{L},\mathcal{E})\cong Ext^{\bullet}_{\mathscr{U}(\mathcal{O}_{X},\mathcal{L})}(\mathcal{O}_{X},\mathcal{E})$. Here, we use the result of Lie-Rinehart cohomology as derived functor for a projective Lie-Rinehart algebra [34, 16] as a local description. We apply this result for the sheaf of logarithmic derivations and for $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras. Furthermore, we express the Lie algebroid cohomology in terms of Čech cohomology by considering a good open cover [33, 36, 2]. As an application of the cohomology groups, we consider abelian Lie algebroid extensions and (first) Chern classes [24, 25] in this general set up. In Section 4, we consider algebraic (analytic) de Rham cohomologies for some free divisors associated with principal ideal sheaves [29] and compute the corresponding logarithmic de Rham cohomologies [4, 27]. We view these cohomologies as Lie algebroid (hyper)cohomology as described in [32, 2, 3].
@@ 2.1. Lie algebroids over algebraic spaces
For an $a$-space $(X,\mathcal{O}_{X})$, the sheaf of $\mathbb{K}_{X}$-linear derivations of $\mathcal{O}_{X}$ is denoted by $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$.
@@
Consider smooth manifolds, complex manifolds, analytic spaces and algebraic varieties, all with their associated structure sheaf, as special $a$-spaces. The tangent sheaf over such an $a$-space $(X,\mathcal{O}_{X})$ is $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, has a standard compatible $\mathbb{K}_{X}$-Lie algebra and a ((((quasi))))coherent $\mathcal{O}_{X}$-module structure [18, 32]. A special $a$-space $(X,\mathcal{O}_{X})$ is said to be non-singular or smooth if its tangent sheaf is a locally free $\mathcal{O}_{X}$-module.
The standard notion of Lie algebroids is defined on a smooth (real) vector bundle over a smooth manifold. The space of global sections of a Lie algebroid forms a Lie-Rinehart algebra, capturing the entire structure. In contrast, for holomorphic (or algebraic) vector bundles over complex manifolds (or algebraic varieties), the space of global sections does not capture the whole information. A sheaf-theoretic approach is necessary for analytic and algebraic geometry, applicable to both singular and non-singular cases.
@@
A homomorphism of Lie algebroids over an $a$-space $(X,\mathcal{O}_{X})$
@@
is a sheaf homomorphism of Lie-Rinehart algebras, i.e.
  • $\phi:\mathcal{L}_{1}\rightarrow\mathcal{L}_{2}$ is an $\mathcal{O}_{X}$-module homomorphism,
  • $\phi:(\mathcal{L}_{1},[\cdot,\cdot]_{1})\rightarrow(\mathcal{L}_{2},[\cdot,\cdot]_{2})$ is a $\mathbb{K}_{X}$-Lie algebra homomorphism,
  • compatibility condition: $\mathfrak{a}_{2}\circ\phi=\mathfrak{a}_{1}$.
@@
The standard Lie algebroid structure on the tangent sheaf of an $a$-space $(X,\mathcal{O}_{X})$ is described by the pair $(\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X}),[\cdot,\cdot]_{c},id)=:\mathcal{T}_{X}$. In particular, when $X$ is a real smooth manifold ((((complex manifold)))) with the structure sheaf $\mathcal{O}_{X}$, the sheaf of smooth ((((holomorphic)))) vector fields $\mathfrak{X}_{X}$ on $X$ is isomorphic to $\mathcal{D}er_{\mathbb{K}_{X}}(\mathcal{O}_{X})$, forming a locally free Lie algebroid of rank equals to dimension of the manifold $X$.
@@ 2.2. Universal enveloping algebroid of a Lie algebroid
From the construction of $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$, it is an associative $\mathbb{K}_{X}$-algebra and $\mathcal{O}_{X}$-bimodule. Moreover, there is a canonical $\mathbb{K}_{X}$-algebra monomorphism $\iota_{X}:\mathcal{O}_{X}\hookrightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ and an $\mathcal{O}_{X}$-linear map $\iota_{\mathcal{L}}:\mathcal{L}\rightarrow\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$. The associative $\mathbb{K}_{X}$-algebra $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$ is generated by $\mathcal{O}_{X}$ and $\iota_{\mathcal{L}}(\mathcal{L})$ satisfy the following identities:
@@ 2.3. $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras
  • $\mathcal{A}$ is equipped with the morphism of $\mathcal{O}_{X}$-modules, the comultiplication $\Delta:\mathcal{A}\rightarrow\mathcal{A}\otimes_{\mathcal{O}_{X}}\mathcal{A}$ and counit $\epsilon:\mathcal{A}\rightarrow\mathcal{O}_{X}$;
@@
For a smooth ((((holomorphic)))) Lie algebroid [2] i.e. for a locally free sheaf of Lie-Rinehart algebras of finite rank over a smooth (complex) manifold, the universal enveloping algebroid is a cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra ((((of finite type)))) [29]. On the other hand, the tangent sheaf over the affine scheme $Spec(\mathbb{A}^{\mathbb{N}})$ ((((where $\mathbb{A}^{\mathbb{N}}:=\mathbb{K}[x_{i}]_{i\in\mathbb{N}})$ [9, 29] and the free Lie algebroid $\mathcal{P}_{X}$ over a smooth manifold $X$ [18] are locally free sheaves of Lie-Rinehart algebras ((((of infinite rank)))). Thus, their universal enveloping algebroids are cocomplete locally graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebras ((((of infinite type)))) [29].
@@ 3. Lie algebroid Cohomology over $a$-spaces
Lie algebroids over an algebraic space (shortly, an $a$-space) and homomorphisms among them are discussed in Section 2.1. These form a category, where kernel and image of a morphism provides new objects in that category. We recall the notion of representations or module of a Lie algebroid in this context. The category of representations of a Lie algebroid has enough injectives (since the category of quasicoherent sheaves of left $\mathcal{O}_{X}$-modules is enough injectives [14]), useful for doing homological algebra.
@@ 3.1. Atiyah algebroid
In particular, when $\mathcal{E}=\mathcal{O}_{X}$ for a non singular $a$-space $X$, we have $\mathcal{A}t(\mathcal{O}_{X})\cong\mathcal{O}_{X}\oplus\mathcal{T}_{X}$, equals to the sheaf differential operators on $X$ of order $\leq 1$ with scalar symbol. The universal enveloping algebroid of the Atiyah algebroid $\mathcal{A}t(\mathcal{O}_{X})$ is the sheaf of twisted differential operators over $X$ [1].
@@
For a $\mathcal{O}$-module $\mathcal{C}$, we can canonically form the following sheaves of graded vector spaces $\wedge^{\bullet}_{\mathcal{O}}$ $\mathcal{C}:=\oplus_{i}\wedge^{i}_{\mathcal{O}}\mathcal{C}$ and $\mathcal{C}^{\otimes^{\bullet}}:=\oplus_{i}$ $\mathcal{C}^{\otimes^{i}}$, where $i\geq 0$ or $i\in\mathbb{N}\cup\{0\}$. These form exterior algebra and tensor algebra of $\mathcal{C}$ with respect to the products $\wedge$ and $\otimes$ over $\mathcal{O}$ respectively.
@@
A quasi-isomorphism of complexes $f^{\bullet}:\mathcal{F}^{\bullet}\rightarrow\mathcal{G}^{\bullet}$ induces an isomorphism $\mathbb{H}^{k}(X,f^{\bullet})$, and if $\mathcal{I}^{\bullet}$ is a complex of injective sheaves then $\mathbb{H}^{k}(X,\mathcal{I}^{\bullet})=H^{k}(\Gamma(X,\mathcal{I}^{\bullet}))$, where $\Gamma(X,\mathcal{I}^{\bullet})$ is the space of global sections of $\mathcal{I}^{\bullet}$.
@@
We use sheaf-theoretic generalizations of the well-known derived functors $Ext$ and $Cotor$ ((((see [34],[19])))), by sheafifying the standard cochain complexes, and subsequently considering the associated hypercohomologies.
@@ 3.2. Chevalley-Eilenberg-de Rham complex and Koszul-Rinehart resolution
where $D_{1},\dots,D_{k+1}\in\mathcal{L}(U)$ and $\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U))$, and $\nabla:~{}U\mapsto\nabla^{U}$ (see the notion in (2)) is the sheafification. Notice that, the differential $d$ (i.e. $d^{2}=0)$ is a $\mathbb{K}_{X}$-linear map, but not an $\mathcal{O}_{X}$-module homomorphism ((((satisfies the Leibniz rule: $d(f~{}\omega)=df\wedge\omega+f~{}d\omega$, for $f\in\mathcal{O}_{X}(U),\omega\in\Omega^{k}_{\mathcal{L}(U)}(\mathcal{E}(U)))$.
@@
which implies $\Omega^{\bullet}_{\mathcal{L}}(\mathcal{E})\cong\Omega^{\bullet}_{\mathcal{L}}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ where $\Omega^{\bullet}_{\mathcal{L}}:=\Omega^{\bullet}_{\mathcal{L}}(\mathcal{O}_{X})$. This complex is equivalent to the standard complex associated with a smooth Lie algebroid [21] or a holomorphic Lie algebroid [32]. In particular, we get the Chevalley-Eilenberg complex when $\mathcal{L}$ is a Lie algebroid over a point; and the de Rham complex $\Omega^{\bullet}_{X}$ when $\mathcal{L}=\mathcal{T}_{X}$ and $\mathcal{E}=\mathcal{O}_{X}$ over a non-singular (or smooth) $a$-space $X$ ((((see Note 2.2)))).
@@
Consider the chain complex of left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-modules, called the Koszul-Rinehart complex:
@@
it is the sheafification of the presheaf of cochain complexes ((((see [34, 40]))))
@@
For every $x\in X$, the space of sections $\mathcal{L}(U_{x})$ is a $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra and a free module over $\mathcal{O}_{X}(U_{x})$. Thus, the Koszul-Rinehart resolution (7) is a locally free resolution of the sheaf of local rings $\mathcal{O}_{X}$. Note that, here the $\mathcal{L}$-module $\mathcal{E}$ can be considered as a $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module.
Here, we consider the two naturally associated presheaf of cochain complexes ((((using notions 3.2 and 3.2))))
@@
One can find the proof for a locally free Lie algebroid $\mathcal{L}$ of finite rank over a Noetherian separated scheme $(X,\mathcal{O}_{X})$ discussed in [3], by using the ideas of [34] on the level of stalks $\mathcal{L}_{x}$ ((((since $\mathcal{L}_{x}$ is a $(\mathbb{K},\mathcal{O}_{X,x})$-Lie-Rinehart algebra, projective (in fact, free) module over $\mathcal{O}_{X,x}$ for all $x\in X)$.
@@
Consider the path algebroid $\mathcal{P}_{X}$ for a non-singular special $a$-space $(X,\mathcal{O}_{X})$ ((((see Note 2.2)))) and its universal enveloping algebroid $\mathbb{D}_{X}$ (see Remark 2.12). Note that, a module over the free Lie algebroid $\mathcal{P}_{X}$ can be viewed as a vector bundle with $\mathcal{T}_{X}$-connection ((((not necessarily flat)))), and a $\mathbb{D}_{X}$-module describes some system of linear PDE on the space of paths (see [18] for details).
@@
Also, note that the whole ideas works for the classical case of Lie algebroids over smooth manifolds (i.e. for smooth Lie algebroids). In [2], the Lie algebroid ((((hyper))))cohomology of holomorphic ((((algebraic)))) Lie algebroids over a complex manifold ((((smooth scheme)))) has expressed by Čech cohomology using a good open cover consisting with connected Stein manifolds ((((affine schemes)))). However, the case associated with analytic spaces is not considered. Additionally, this result can be extended to the quasicoherent $\mathcal{O}_{X}$-module cases.